How many combinations are there for a lock with four entries, where each entry can be 0-9, a-z, or A-Z?
Answer #1

Total possible entries for one place = 10 + 26 + 26 = 62

So, total combinations = 62 * 62 * 62 * 62 = 14776336

Hence, total 14776336 locks with four entires are possible.

Please comment if any doubt. Thank you.